A 0.2 kg cannon ball is fired at an upward angle of 45° from the top of a 165 m cliff with a speed of 175 m/s. (A) Using conservation of energy, find the speed it has when it strikes the ground below. Suppose the cannon ball lands in a soft, muddy field. The force of the field is 75 N on the projectile. (B) How deep does it penetrate into the ground?

--------------------------------------------------------------------------------

How much energy does it take to lift a 5 lb brick to a height of 1 meter?

One barrel of oil has a chemical potential energy of approximately 6 x 109 J. How high would the energy from a barrel of oil lift the same 5 lb brick?

How many times per second would you have to lift the brick to get 100 W of power?
Answer #1

m = mass of cannonball = 0.2 kg

\theta = angle from horizontal = 45

h = height = 165 m

V = speed at top = 175 m/s

V' = speed at bottom

A)

using conservation of energy

Potential energy at Top + Kinetic energy at top = Kinetic energy at bottom + Potential energy at bottom

mgh + (0.5) m V2 = (0.5) m V'2 + mg (0)

V'2 = V2 + 2gh

V'2 = 1752 + 2 x 9.8 x 165

V' = 184.01 m/s

B)

a = retardation caused inside the ground = - F/m = - 75 / 0.2 = - 375 m/s2

Vf = final speed = 0

d = displacement inside the ground

using the equation

Vf2 = V'2 + 2 a d

02 = 184.012 + 2 (-375) d

d = 45.2 m

m = 5 lb = 2.27 kg

h = 1 m

energy = potential energy gained = mgh = 2.27 x 1 x 9.8 = 22.25 J